OpenStudy (anonymous):

b and c are in a proportional relationship. When b=4, c=16. What is the value of b when c=24?

OpenStudy (ipwnbunnies):

Proportional means both b and c should get bigger and smaller together. So: b/c = b/c The first b/c is for the given values. You need to solve for b on the right side

OpenStudy (anonymous):

so i would just do 16 times 24 divide by 4 right or no

OpenStudy (abdullah1995):

@maddieleee that is a really disgusting and rude statement.

OpenStudy (one098):

@abdullah1995 she's a troll. just ignore and report. she's just a disgusting person in every way.

OpenStudy (abdullah1995):

\[4/16 = x/24\]

OpenStudy (abdullah1995):

@One098 i guess that's the best approach

OpenStudy (one098):

yeah

OpenStudy (anonymous):

thank you all of you was great help